OpenStudy (anonymous):

Evaluate f(x) = x² + 1 for f(10). a.2√2 b.101 c.21

OpenStudy (anonymous):

b

OpenStudy (callisto):

f(10) = 10² + 1 =100+1=101=B

OpenStudy (anonymous):

lol, 101 sry, many faster than me

OpenStudy (callisto):

no problem lol